# Effect of primary scalar hair on black hole’s strong lensing in Beyond Horndeski Gravity

### v1

@@ V.2 Effects of the primary hair on the higher-order images
Thus, it is clear that the tertiary image is not a useful tool, as it is thin and nearly attached to the inner black hole shadow. Therefore, it is not observed as the photon ring in GRMHD simulations. As discussed in Ref. [103], this evidence suggests that higher-order images (black hole shadow subrings) are not of significant astrophysical application, similar to black hole subrings.
@@ VI Summary and conclusions
A key aspect of our study was the formulation of the deflection angle in the strong lensing regime. Utilizing the Virbhadra-Ellis lens equation and Bozza’s analytical approach, we demonstrated how the primary scalar hair parameter influences the bending of light rays around the black hole. Our results show that the presence of primary scalar hair alters the location and properties of the photon sphere, leading to significant modifications in the lensing observables. Notably, we found that black holes with negative scalar hair parameter exhibit a larger deflection angle compared to their general relativistic counterparts, while those with positive parameter display reduced deflection effects. This observation directly impacts the formation of relativistic images and the Einstein rings observed in black hole imaging.
@@
Finally, an important conclusion of our study is the local instability of black hole solutions satisfying the derived constraints. As previously noted, models with primary scalar hair within the range $-9.097\leq c\leq 1.319$ exhibit instabilities, raising concerns about their astrophysical relevance. This suggests that while observational constraints favor specific values of the scalar hair parameter, the physical viability of these models requires further investigation, particularly in the context of dynamical stability and perturbative analyses.
In summary, our work presents a detailed examination of strong gravitational lensing in Beyond Horndeski Gravity and provides new constraints on the primary scalar hair parameter using observational data from EHT. These findings contribute to the ongoing effort to test and refine alternative theories of gravity through astrophysical observations, paving the way for future studies on black hole phenomenology beyond general relativity.

### v2

@@ V.2 Effects of the primary hair on the higher-order images
Thus, the tertiary image appears as a very thin structure, located very close to the inner black hole shadow, making its separation challenging in practical observations. Moreover, current VLBI angular resolutions, flux limitations, and $(u,v)$ coverage further hinder the detectability of such higher-order images, as they are effectively unresolved and faint in GRMHD simulations. As discussed in Ref. [114], these observational constraints imply that higher-order images (black hole shadow subrings) are unlikely to provide significant astrophysical information with current instrumentation, analogous to the challenges faced in resolving black hole subrings.
@@ VI Summary and conclusions
A key aspect of our study was the formulation of the deflection angle in the strong lensing regime. Utilizing the Virbhadra-Ellis lens equation and Bozza’s analytical approach, we demonstrated how the primary scalar hair parameter influences the bending of light rays around the black hole. Our results show that the presence of primary scalar hair alters the location and properties of the photon sphere, leading to significant modifications in the lensing observables. Notably, we found that black holes with negative scalar hair parameter exhibit a larger deflection angle compared to their general relativistic counterparts, while those with positive parameter display reduced deflection effects. This observation directly impacts the formation of relativistic images and the RERs observed in black hole imaging.
@@
Finally, an important conclusion of our study is the local instability of black hole solutions satisfying the derived constraints. As previously noted, models with primary scalar hair within the range $9.097\leq c\leq 1.319$ exhibit instabilities, raising concerns about their astrophysical relevance. This suggests that while observational constraints favor specific values of the scalar hair parameter, the physical viability of these models requires further investigation, particularly in the context of dynamical stability and perturbative analyses.
In summary, this work provides a comprehensive analysis of strong gravitational lensing in Beyond Horndeski gravity, deriving new observational constraints on the primary scalar hair parameter using the EHT data. These results advance the testing and refinement of alternative gravity theories through astrophysical measurements, offering valuable insights into black hole properties beyond general relativity. Future research may build upon this foundation by investigating strong lensing effects in a wider range of black hole solutions with scalar hair recently explored in the literature [119, 120]. Although such studies pose significant computational challenges, they hold promise for deepening our understanding of the physical viability and thermodynamic stability of these extended scalar-tensor models [121, 122].
Finally, we note that the static black hole solutions studied here, as well as similar configurations, can in principle be extended to rotating counterparts via the Newman–Janis algorithm [123]. The strong gravitational lensing analysis could then be applied to these rotating geometries using the formalism developed in Ref. [59]. Such an extension would offer a more realistic framework for astrophysical applications and is left for future work.